OpenStudy (anonymous):

Need help finding solution for non-homogeneous initial value problem. --See below--

OpenStudy (anonymous):

y'' - 2y' + y = te^t + 4 iconditions; y(0) = 0; y'(0) = 0 I don't even know where to start :( should i find general solution for homogeneous side and then find y(t) with deriv = (te^t + 4)!!

OpenStudy (jamesj):

So three steps 1. Find the homogeneous solutions, y1 and y2 to y'' - 2y' + y = 0 2. Find the particular solution yp to the inhomogeneous equation 3. Write down the general solution y = c1.y1 + c2.y2 + yp and substitute the initial solutions to find the final solution

OpenStudy (jamesj):

So now, I hope step 1 is easy.

OpenStudy (jamesj):

...easy for you. yes?

OpenStudy (anonymous):

yes, then solve separately and add solutions?

OpenStudy (jamesj):

yes. Now step 2. Easiest thing here is the variation of unknown parameters.

OpenStudy (jamesj):

Your text book should talk about this

OpenStudy (jamesj):

The general form I'd use to substitute with this method is yp = Ate^t + Be^t + C. Substitute into the equation and you'll be able to solve for A, B and C.

OpenStudy (anonymous):

Thanks. I wasn't sure which method to pursue, I'll look up the variation of parameters section.